OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? y = x + 1 y = 1 y = 2x + 1 y = 5x + 8

OpenStudy (anonymous):

the answer is a because y = x + 1 plug it in 5 = 4 + 1, 9 = 8 + 1

OpenStudy (anonymous):

witch one i sthat

OpenStudy (anonymous):

the first one A

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

Yup!